OpenStudy (dukelaxfan):

The pool maintenance man forgot to bring his logarithmic charts, and he needs to raise the amount of hydronium ions, t, in the pool by 0.50. To do this, he can use the graph you created. Use your graph to find the pH level if the amount of hydronium ions is raised to 0.50. Then, convert the logarithmic function into an exponential function using y for the pH.

OpenStudy (dukelaxfan):

@ganeshie8 @whpalmer4 @mathstudent55

OpenStudy (dukelaxfan):

Please help

OpenStudy (dukelaxfan):

@mathmale

OpenStudy (aravindg):

Do you know the equation for pH?

OpenStudy (dukelaxfan):

If you ever swam in a pool and your eyes began to sting and turn red, you felt the effects on an incorrect pH level. pH measures the concentration of hydronium ions and can be modeled by the function p(t) = −log10t. The variable t represents the amount of hydronium ions; p(t) gives the resulting pH level. Water at 25 degrees Celsius has a pH of 7. Anything that has a pH less than 7 is called acidic, a pH above 7 is basic, or alkaline. Seawater has a pH just more than 8, whereas lemonade has a pH of approximately 3.

OpenStudy (aravindg):

Okay so in our question, concentration of ions is increased by 1/2. Can you form the equation for that from basic equation of pH?

OpenStudy (dukelaxfan):

would the base be 0.5?

OpenStudy (aravindg):

Never! pH is always done with base 10. You only have to change the concentration part!

OpenStudy (dukelaxfan):

ok, sorry, im just a bit slow with this

OpenStudy (aravindg):

It's okay. Take your time :)

OpenStudy (dukelaxfan):

p(h)=log(10)(0.5)

OpenStudy (dukelaxfan):

\[\log_{10} (0.5)\]

OpenStudy (aravindg):

If the original equation is p(t) = −log10t. Wouldn't it be p(t) = −log10(t+0.5) ??

OpenStudy (dukelaxfan):

that makes sense, like i said im a little slow with this so itll be a bit hard for me

OpenStudy (aravindg):

Okay. Can you try it with that input?

OpenStudy (dukelaxfan):

i dont know what you mean.

OpenStudy (dukelaxfan):

Graph it?

OpenStudy (dukelaxfan):

I can do that.

OpenStudy (aravindg):

Let's see that since the question asks you to.

OpenStudy (aravindg):

That format doesnt work on my pc.

OpenStudy (dukelaxfan):

o, let me try another, sorry about that

OpenStudy (dukelaxfan):

does that work?

OpenStudy (aravindg):

Good. Sorry got bsy.

OpenStudy (dukelaxfan):

no problem

OpenStudy (aravindg):

Now you need to convert it into exponential function For that write p(t) as y then take 10^y from the equation.

OpenStudy (dukelaxfan):

y=-log(t+0.5)?

OpenStudy (aravindg):

yes take 10^y from that.

OpenStudy (dukelaxfan):

y-10^y=-log(t+0.5)?

OpenStudy (aravindg):

Nope. First -log A=log 1/A Use that

OpenStudy (dukelaxfan):

what does the A stand for, (t+0.5)?

OpenStudy (aravindg):

Yup

OpenStudy (dukelaxfan):

ok

OpenStudy (dukelaxfan):

\[-\log_{10} (t+0.5)=\log(1)/(t+0.5) \]

OpenStudy (dukelaxfan):

thats what i interpreted

OpenStudy (aravindg):

This equals y So 10^y=log(1/t+0.5)) This is the expression in exponential form.

OpenStudy (dukelaxfan):

ok, thank you so much. I really appreciate it.

OpenStudy (aravindg):

yw :)

OpenStudy (aravindg):

Wait u shoudn't write log in the final equation! Sorry abt that!

OpenStudy (dukelaxfan):

so 10^y=(1/(t+0.5))

OpenStudy (aravindg):

Yep.

OpenStudy (dukelaxfan):

ok thx

OpenStudy (dukelaxfan):

im gonna close the question now

OpenStudy (dukelaxfan):

ok?

OpenStudy (aravindg):

Sure! This indeed took some time isn't it? :)

OpenStudy (dukelaxfan):

yes